OpenStudy (shadowcain):

Which equations are equivalent to the given equation? 0.4y = 10 Choose all answers that are correct. A. 0.4y = 10 / .04 B. 0.4 / 0.4 = 10 / 0.4 C. 0.4 X 2.5 = 10 X 2.5 D. 0.4y / 0.4 = 10 X 0.4

OpenStudy (freckles):

I think there is a type-o some of the equations don't even have y in it

OpenStudy (madhu.mukherjee.946):

yes

OpenStudy (freckles):

To solve something like ay=b for y you need to divide both sides by a

OpenStudy (madhu.mukherjee.946):

i think if b had y in it ,it would have been the correct answer

OpenStudy (madhu.mukherjee.946):

so i'll go for B

OpenStudy (shadowcain):

A. 0.4y = 10 / .04 y = 625 B. 0.4y / 0.4 = 10 / 0.4 C. 0.4y X 2.5 = 10 X 2.5 D. 0.4y / 0.4 = 10 X 0.4 Sorry I typed it wrong

OpenStudy (madhu.mukherjee.946):

y=625 meaning????is that the answer

OpenStudy (shadowcain):

for A. y= 625 I think. I was just trying to figure them all out to narrow them down.

OpenStudy (madhu.mukherjee.946):

not 625 its 62.5

OpenStudy (madhu.mukherjee.946):

i told you its option b

OpenStudy (shadowcain):

Is that the only one? Because it says to choose all correct answers.

OpenStudy (madhu.mukherjee.946):

also option C

OpenStudy (shadowcain):

Thanks for your help

OpenStudy (madhu.mukherjee.946):

welcome